# Safe Learning for Contact-Rich Robot Tasks: A Survey from Classical Learning-Based Methods to Safe Foundation Models

### v1

@@ 3.8.3 Dual-Space Safety Enforcement
While enforcing safety solely in either task space or joint space can already reduce risk, each perspective has blind spots. Task-space constraints may miss internal joint-limit or actuator issues, whereas purely joint-space safeguards may overlook subtle but harmful contact patterns at the end-effector.
For this reason, some frameworks combine safety mechanisms in both the task space and the joint space to provide a more comprehensive safety net. In these approaches, the robot monitors and constrains its behavior at the end-effector level and at the actuator level simultaneously. For example, a hierarchical safe control scheme might use a Cartesian impedance controller to ensure compliant, force-limited contact with the environment, while a parallel joint-space module detects unexpected collisions or excessive torques and reacts by altering joint commands to dissipate force safely [66]. This dual-layer design guarantees that even if an unforeseen contact occurs that was not accounted for by the task-space planner, the joint-space safety reflex can kick in to prevent damage.
Another representative work uses a two-loop control architecture for a robot with flexible joints. An outer loop regulates the contact force in task space to track a desired force trajectory, and an inner loop stabilizes the joint motion [56]. Through Lyapunov-based analysis, it is proven that both the force error and the joint deviations remain bounded, illustrating theoretical safety in both domains. In learning-based settings, ensuring stability in the joint space can indirectly enforce task-space safety. For instance, a model-based actor–critic method confines the joint controller’s behavior within stability limits using an analytical model of the robot and environment stiffness. As a result, the end-effector never pushes beyond safe force limits because the joint controller won’t allow an unsafe regime [88]. Even in dynamic locomotion tasks, a similar philosophy is applied.
One study on robotic hopping ensured the policy respected joint torque and angle limits at all times, which inherently kept the robot’s jumping motions stable and prevented falls or hardware stress during ground contact [101]. By addressing safety at multiple levels, these approaches leverage the strengths of each domain. The task-space layer handles safe interaction with objects and surfaces, while the joint-space layer ensures the robot’s own mechanics remain within secure operating conditions. This redundancy provides robust safety even under varying uncertainties or if one layer’s assumptions are violated.
@@ 3.8.4 Policy Spaces
In addition to physical control spaces, researchers have also explored safety in the policy or decision-making space. That is, integrating safety considerations into the robot’s learning and planning algorithms themselves. Rather than hard limits on forces or torques, these methods introduce abstract constraints or safety-centric objectives that guide the policy toward safe behavior.
One common approach is to formulate the problem as a CMDP, where certain state-action pairs or outcomes are labeled as unsafe with an associated cost. The reinforcement learning agent must then optimize its policy while satisfying a constraint on this unsafe cost, often via Lagrange multiplier techniques, to ensure it never exceeds a tolerated risk level [33]. In essence, the policy learns to avoid any behavior that would trigger the unsafe cost, adhering to safety requirements by design.
Another strategy is to learn a safety critic or value function that predicts the long-term risk of states and actions. For example, an algorithm might train a feasibility value function that estimates the probability of constraint violation for a given situation. The policy is then constrained or adjusted using this knowledge, selecting only actions that keep the robot in low-risk regions [45]. Along these lines, a recovery RL framework can learn a special recovery policy that intervenes whenever the agent is near an unsafe state. If the safety critic signals high risk for the agent’s intended action, control is handed over to the recovery policy, which drives the system back to a safe region before returning control to the main policy [91]. In high-dimensional problems, like vision-based manipulation, some works first learn a compact latent state representation and define safety in that latent space. For instance, a variational autoencoder might encode images of the robot workspace into a latent vector, and a learned barrier function on this latent space separates safe and unsafe states. The RL policy is then trained with an additional constraint to remain on the safe side of the barrier, even without explicit physical measurements [65].
There are also human-in-the-loop approaches where safety is addressed at a policy level through human feedback and oversight. In such frameworks, the robot’s exploration is guided by human evaluators or pre-defined safety rules. The robot might query a human or use a human-provided value function to ensure its learned policy aligns with human safety preferences and avoids dangerous actions during training [54].
Lastly, recent generative approaches ensure safe decision sequences by design. For example, diffusion-based policy models can be guided to prefer smooth, low-uncertainty action sequences, which reduces the chance of sudden erratic behavior that could be unsafe [42]. By incorporating safety directly into the decision-making process, these approaches aim to prevent unsafe commands from ever being chosen, complementing the lower-level safety controllers. This policy-space perspective ensures that the robot’s behavior remains within acceptable risk bounds from the planning stage through to execution, adding an extra layer of safety on top of physical enforcement.
@@ 4 Key Challenges
Despite rapid progress in learning-based methods for contact-rich robotic manipulation, several key challenges remain before these approaches can be deployed safely and reliably at scale. This section outlines major open problems, with particular attention to how emerging VLM/VLA-based methods both exacerbate and help address these issues.
@@ 4.1 Fragile Safety and Stability Guarantees
Learning-based controllers, including those using deep networks and VLAs, often lack formal guarantees of passivity, stability, and constraint satisfaction under hard contacts [zhang2025passivity, spyrakos2020passivity, 75], which can lead to unsafe behaviors and hardware damage. While passivity-centric and energy-constrained formulations offer promising directions, integrating them seamlessly with large, expressive policies remains technically challenging, especially when contact conditions change rapidly or unpredictably.
@@ 4.4 Safe Exploration Under Physical Risk
Exploration remains the most safety-critical phase: naive probing causes jamming, surface damage, or tool wear in insertion, polishing, and assembly contexts [yuan2022safe, zhang2024srl]. Existing constrained and risk-sensitive reinforcement learning frameworks provide expectation or probabilistic bounds, yet exhibit weakness under distribution shift. Shielding and action-projection (barrier / reachability / model-predictive) layers offer hard online guarantees, but typically depend on conservative uncertainty sets, limiting data efficiency.
@@ 4.5 Safe Execution and Performance Trade-offs
Ensuring sustained constraint satisfaction post-training (safe execution) requires balancing responsiveness, task efficiency, and physical margin. Variable impedance or admittance policies learned from demonstrations or reinforcement signals modulate compliance adaptively [28, 25, zhang2024srl], yet stability margins under rapid re-tuning are seldom quantified. Runtime safety filters can erode performance through excessive intervention, while soft filtering strategies risk cumulative force violations. There is insufficient methodology for joint optimization over (i) violation probability; (ii) force/energy envelopes; and (iii) efficiency.
@@ 4.6 Benchmark and Dataset Limitations
Existing simulation suites and benchmarks provide limited coverage of deformable objects, multi‑stage processes, tool‑supported tasks, and tasks performed in close collaboration with humans. [towers2024gymnasium, geng2025roboverse, gu2025robust]. Public datasets rarely include (i) high-frequency wrench traces aligned with failure annotations; (ii) rich “near-miss” episodes essential for shaping recovery policies; or (iii) semantic safety labels (forbidden regions, fragile surfaces). This scarcity impedes reproducible comparison of violation metrics, recovery latency, and generalization. Emerging safety-oriented RL frameworks and robust benchmarks partially address constrained exploration but lack standardized contact-force evaluation protocols [152, raffin2021stable].
@@ 4.8 Human-Robot-Collaboration Safety
Human proximity or direct physical collaboration introduces soft constraints (comfort, predictability) alongside hard limits (force, speed). For contact-rich tasks near or with humans, operators must be able to anticipate robot behavior and intervene effectively, yet current learning-based controllers often lack transparency and interpretable failure modes. Existing safe learning pipelines predominantly encode physical thresholds while neglecting temporal predictability or trust metrics central to acceptance [346, 253, hanna2020towards]. Adaptive safety frameworks that integrate ergonomic and standard-driven constraints (e.g., force envelopes in ISO/TS 15066 of [ISO15066]) with learning-enabled modulation are limited. Advanced models, including VLAs, can in principle explain intentions or request help, but designing interaction protocols, interfaces, and explanations that are both faithful and useful in high-contact scenarios remains an open challenge.
@@ 5.1 Standardized Safe-contact Benchmarks
Current suites under-represent multi-stage forceful interaction, deformables, and human-adjacent tasks [towers2024gymnasium, geng2025roboverse, gu2025robust, yuan2022safe]. Future benchmarks for safe learning in contact-rich robotics should move beyond isolated manipulation tasks to suites that stress high-force interactions, tight clearances, and diverse materials and fixtures. A key opportunity is to design benchmarks where vision–language models (VLMs) and vision–language–action (VLA) models must interpret natural language task descriptions and safety rules, then execute contact-rich skills under explicit constraints on forces, slippage, and failure recovery.
@@ 5.4 Multimodal Semantic Grounding
VLM/VLA models enable symbolic specification of scene understanding, semantic reasoning[33, 330, 10705419], yet robust grounding remains fragile under occlusion and clutter. A promising direction is layered grounding, which proceeds from scene parsing to affordance and constraint extraction, followed by formal constraint compilation (e.g., CBF/MPC sets), and ultimately runtime conformance validation. In addition, safety‑aware prompting and structured action frameworks can reduce the risk of hallucinated unsafe plans.
@@ 5.5 Human-Centric and Lifecycle-aware Safety
Physical safety envelopes, such as limits on force and speed, are necessary but not sufficient for collaborative adoption [ajoudani2018progress]. Human-centered safe learning frameworks can benefit from language-conditioned policies that understand instructions, preferences, and corrections expressed in natural language, enabling non-expert users to shape safe behavior even in contact-rich collaboration. Over the system lifecycle, advanced models such as VLAs can support continual learning and self-diagnostics, where robots explain their contact strategies, flag emerging risks, and update policies while preserving guarantees about human safety and equipment integrity.
5.6 Scalable Verification for Learning-based Controllers
@@ 5.8 Open Problems
Key unresolved issues include: (i) certified generalization under simultaneous geometric and material shifts [das2025towards]; (ii) synthesis of minimally conservative multi-contact safe sets; (iii) semantic-to-physical constraint robustness; and (iv) lifecycle adaptation with formal transient bounds. Addressing these will hinge on tighter integration of analytical safety structures with data-driven representations and foundation models.

### v2

@@ 3.8.3 Dual-Space Safety Enforcement
Exclusive reliance on either task or joint space constraints can create safety blind spots. Task space control may inadvertently command infeasible joint configurations, while joint-space safety may overlook excessive contact pressures at the end-effector. Consequently, hybrid architectures have emerged to enforce safety concurrently across both domains.
Hierarchical control frameworks typically employ a task-space loop to manage environmental interaction forces, coupled with an inner joint-space loop to handle redundancy resolution and torque limits. A Cartesian impedance controller regulates contact stiffness, while a secondary joint-space reflex module monitors for collisions or internal limits, overriding commands to dissipate energy if discrepancies arise [399, 319]. Theoretical guarantees for such dual-loop architectures have been established via Lyapunov analysis, proving that errors in both force tracking and joint deviation remain bounded [359, 215]. Additionally, in [179], the constrained manifold is imposed in both joint and Cartesian space to ensure safety.
Ensuring stability in the joint space can indirectly enforce task-space safety [189]. In [386], a model-based actor–critic method confines the joint controller’s behavior within stability limits using an analytical model of the robot and environment stiffness. As a result, the end-effector also never pushes beyond safe force limits. Similarly, in dynamic locomotion tasks like hopping, strictly respecting joint torque and angle limits is shown to inherently stabilize the task-space center-of-mass trajectory, preventing falls [275].
@@ 3.8.4 Policy Spaces
Abstracting away from real-time control, safety is also enforced at the decision-making level. Policy-space methods integrate safety considerations directly into the learning objective or planning horizon, aiming to preemptively avoid states that would necessitate low-level emergency intervention.
A standard formalism for this is the CMDP, where the agent optimizes a reward signal subject to a cost threshold associated with unsafe states [4, 66, 327, 356]. Lagrangian relaxation is commonly used to dynamically weight this cost, training the policy to satisfy safety constraints asymptotically [367]. Alternatively, safety can be encoded into the value function itself[207, 294]. Safety critics or feasibility value functions estimate the long-term probability of constraint violation, guiding the policy to select actions that maximize the margin of safety [94]. To handle potential failures, [301] introduces a separate safety policy that intervenes when the agent approaches a high-risk region, steering the system back to a safe state before returning control to the task policy.
Implementing such boundaries becomes non-trivial when the state space is high-dimensional or unstructured. In high-dimensional visual manipulation tasks, safety barriers are often learned in latent spaces, allowing the agent to distinguish between safe and unsafe visual manifolds without explicit state estimation [366, 294, 330]. Furthermore, when safety constraints are difficult to formalize mathematically, Human-in-the-Loop paradigms incorporate human oversight to shape the policy’s risk sensitivity, utilizing human feedback and intervention to define safety boundaries during the exploration phase [89]. Beyond defining constraints and boundaries, generative models such as diffusion policies have been employed to synthesize action sequences that prioritize smoothness and low uncertainty, thereby reducing the stochasticity that often leads to unsafe behaviors [46, 95, 375].
@@ 4 Key Challenges
Despite rapid progress in learning-based methods for contact-rich robotic manipulation, several key challenges remain before these approaches can be deployed safely and reliably at scale. This section outlines major open problems, with particular attention to how emerging VLM/VLA-based methods both amplify and help address these issues.
@@ 4.1 Fragile Safety and Stability Guarantees
Learning-based controllers, including those using deep networks and VLAs, often lack formal guarantees of passivity, stability, and constraint satisfaction under hard contacts [373, 283, 131], which can lead to unsafe behaviors and hardware damage. While passivity-centric and energy-constrained formulations offer promising directions, integrating them seamlessly with large, expressive policies remains technically challenging, especially when contact conditions change rapidly or unpredictably.
@@ 4.4 Safe Exploration Under Physical Risk
Exploration remains the most safety-critical phase: naive probing causes jamming, surface damage, or tool wear in insertion, polishing, and assembly contexts [362, 374]. Existing constrained and risk-sensitive reinforcement learning frameworks provide expectation or probabilistic bounds, yet exhibit weakness under distribution shift. Shielding and action-projection (barrier, reachability, model-predictive) layers offer hard online guarantees, but typically depend on conservative uncertainty sets, limiting data efficiency.
@@ 4.5 Safe Execution and Performance Trade-offs
Ensuring sustained constraint satisfaction post-training (safe execution) requires balancing responsiveness, task efficiency, and physical margin. Variable impedance or admittance policies learned from demonstrations or reinforcement signals modulate compliance adaptively [169, 351, 374], yet stability margins under rapid re-tuning are seldom quantified. Runtime safety filters can erode performance through excessive intervention, while soft filtering strategies risk cumulative force violations. There is insufficient methodology for joint optimization over (i) violation probability; (ii) force/energy envelopes; and (iii) efficiency.
@@ 4.6 Benchmark and Dataset Limitations
Existing simulation suites and benchmarks provide limited coverage of deformable objects, multi‑stage processes, tool‑supported tasks, and tasks performed in close collaboration with humans. [308, 82, 91]. Public datasets rarely include (i) high-frequency wrench traces aligned with failure annotations; (ii) rich “near-miss” episodes essential for shaping recovery policies; or (iii) semantic safety labels (forbidden regions, fragile surfaces). This scarcity impedes reproducible comparison of violation metrics, recovery latency, and generalization. Emerging safety-oriented RL frameworks and robust benchmarks partially address constrained exploration but lack standardized contact-force evaluation protocols [123, 243].
@@ 4.8 Human-Robot-Collaboration Safety
Human proximity or direct physical collaboration introduces soft constraints (comfort, predictability) alongside hard limits (force, speed). For contact-rich tasks near or with humans, operators must be able to anticipate robot behavior and intervene effectively, yet current learning-based controllers often lack transparency and interpretable failure modes. Existing safe learning pipelines predominantly encode physical thresholds while neglecting temporal predictability or trust metrics central to acceptance [199, 263, 99]. Adaptive safety frameworks that integrate ergonomic and standard-driven constraints (e.g., force envelopes in ISO/TS 15066 of [119]) with learning-enabled modulation are limited. Advanced models, including VLAs, can in principle explain intentions or request help, but designing interaction protocols, interfaces, and explanations that are both faithful and useful in high-contact scenarios remains an open challenge.
@@ 5.1 Standardized Safe-contact Benchmarks
Current suites under-represent multi-stage forceful interaction, deformables, and human-adjacent tasks [308, 82, 91, 362]. Future benchmarks for safe learning in contact-rich robotics should move beyond isolated manipulation tasks to suites that stress high-force interactions, tight clearances, and diverse materials and fixtures. A key opportunity is to design benchmarks where VLM/VLA models must interpret natural language task descriptions and safety rules, then execute contact-rich skills under explicit constraints on forces, slippage, and failure recovery.
@@ 5.4 Multimodal Semantic Grounding
VLM/VLA models enable symbolic specification of scene understanding, semantic reasoning[367, 307, 388], yet robust grounding remains fragile under occlusion and clutter. A promising direction is layered grounding, which proceeds from scene parsing to affordance and constraint extraction, followed by formal constraint compilation (e.g., CBF/MPC sets), and ultimately runtime conformance validation. In addition, safety‑aware prompting and structured action frameworks can reduce the risk of hallucinated unsafe  [298].
@@ 5.5 Human-Centric and Lifecycle-aware Safety
Physical safety envelopes, such as limits on force and speed, are necessary but not sufficient for collaborative adoption [6]. Human-centered safe learning frameworks can benefit from language-conditioned policies that understand instructions, preferences, and corrections expressed in natural language, enabling non-expert users to shape safe behavior even in contact-rich collaboration. Over the system lifecycle, advanced models such as VLAs can support continual learning and self-diagnostics, where robots explain their contact strategies, flag emerging risks, and update policies while preserving guarantees about human safety and equipment integrity.
5.6 Scalable Verification for Learned Controllers
@@ 5.8 Open Problems
Key unresolved issues include: (i) certified generalization under simultaneous geometric and material shifts [62]; (ii) synthesis of minimally conservative multi-contact safe sets; (iii) semantic-to-physical constraint robustness; and (iv) lifecycle adaptation with formal transient bounds. Addressing these will hinge on tighter integration of analytical safety structures with data-driven representations and foundation models.